OpenStudy (eric_d):

find inverse h(x)=x+2/x-2 , x not equal to 2 @hartnn

hartnn (hartnn):

let y = (x+2)/(x-2) our aim is to isolate 'x' start by multiplying x-2 on both sides

OpenStudy (eric_d):

ok

OpenStudy (eric_d):

after that

OpenStudy (eric_d):

y(x-2)=x+2

hartnn (hartnn):

distribute 'y' on left

OpenStudy (eric_d):

how..

hartnn (hartnn):

\(\large ☻(☺+♥) = ☻☺+☻♥ \\ \text {this is distributing}\)

OpenStudy (anonymous):

\[y=\frac{ x+2 }{ x-2 }\]\[y->x , x->y\]\[x=\frac{ y+2 }{ y-2 }\]\[xy-2x=y+2\]\[y(x-1)=2x+2\]\[y=\frac{ 2(x+1) }{ x-1 }\]

hartnn (hartnn):

cody, please don't give away the steps, let eric try it on his own.

hartnn (hartnn):

y(x-2)=x+2 could you now distribute the 'y' ?

OpenStudy (eric_d):

Alrite, understood Thank you

OpenStudy (eric_d):

Thanks @hartnn for pointing out the key word...